OpenStudy (anonymous):

Discrete math: Use induction to show that for all positive integers n (a) 13 + 23 + 33 + . . . + n 3 = (n(n + 1)/2)2 . (b) 1 · 1! + 2 · 2! + . . . + n · n! = (n + 1)! − 1 (c) if n > 6, then 3n < n! I did some work for a, and b but im stuck: a--> assume for 1 1^3 = (1(1+1)/2)^2 which is correct Hypothesis: 1^3+2^3+...+k^3= (k(k+1)/2)^2 Induction step: 1^3+2^3+...+k^3+(k^3+1)= (k(k+1)/2)^2 + (k^3 + 1) I started with the RHS and got this: (k^4 + 6k^3 + k^2)/4 + 1 but can't seem to get back to the original rule for b: both Hypothesis and n(1) work

OpenStudy (anonymous):

works, so the Induction step left me with 2(k+1)! +k which I can't seem to bring back to the original rule too, any help? :/

OpenStudy (perl):

you would like help with part a) ?

OpenStudy (anonymous):

with all of them, I can't find a way to simplify them I tried to do them all but I get stuck at the induction step somewhere in my simplification. Both A and B my simplification is wrong somewhere

OpenStudy (perl):

$$\Large \rm { Assume~ that:\\ 1^3+2^3+...+k^3= (k(k+1)/2)^2 \\ then~ it~ follows~ \\ 1^3+2^3+...+k^3 + (k+1)^3 \\ = \color{blue}{ (k(k+1)/2)^2} + (k+1)^3 }$$

OpenStudy (perl):

do you see the substitution there

OpenStudy (anonymous):

Yeah totally see it, but my simplification is wrong somewhere, i mean it shouldn't be hard those are just quadratics

OpenStudy (anonymous):

oh wait... shouldnt it be k^3+1 ?

OpenStudy (anonymous):

because k^3 and the next number would be k^3 +1?

OpenStudy (perl):

no, the next number is (k+1)^3, after k^3

OpenStudy (perl):

For example: if k = 100 , then k+1 = 101 k^3 = 100 ^3 (k+1)^3 = (100+1)^3 = 101^3

OpenStudy (perl):

we are cubing numbers, not simplying adding 1

OpenStudy (anonymous):

Ohhh I see, k^3 the next number would the the following k exponent 3, so (k+1)^3. My bad you are right

OpenStudy (anonymous):

then you end up with (k^4 + 14k^3 + 13k^2 +3k + 4 )/4, you can take out 4/4 to get your one, but seems hard to get back to where we were

OpenStudy (perl):

$$ \Large \rm { Assume~ that:\\ 1^3+2^3+...+k^3= \left( \frac{k(k+1)}{2} \right) ^2 \\ then~ it~ follows~ \\ 1^3+2^3+...+k^3 + (k+1)^3 \\ = \color{blue}{ \left( \frac{k(k+1)}{2} \right) ^2} + (k+1)^3 \\ = \frac{k^2(k+1)^2}{4} + (k+1)^3 \\ = \frac{k^2(k+1)^2}{4} +\frac{4\cdot (k+1)^3}{4} \\ = \frac{k^2(k+1)^2+4\cdot (k+1)^3}{4} \\ = \frac{(k+1)^2\cdot [k^2 + 4(k+1)]}{4} } $$

OpenStudy (anonymous):

oh I see you didn't bother expandiing

OpenStudy (anonymous):

even this though doesn't represent the initial rule though does it ?

OpenStudy (perl):

we can factor the quadratic

OpenStudy (perl):

$$ \Large \rm { Assume~ that:\\ 1^3+2^3+...+k^3= \left( \frac{k(k+1)}{2} \right) ^2 \\ then~ it~ follows~ \\ 1^3+2^3+...+k^3 + (k+1)^3 \\ = \color{blue}{ \left( \frac{k(k+1)}{2} \right) ^2} + (k+1)^3 \\ = \frac{k^2(k+1)^2}{4} + (k+1)^3 \\ = \frac{k^2(k+1)^2}{4} +\frac{4\cdot (k+1)^3}{4} \\ = \frac{k^2(k+1)^2+4\cdot (k+1)^3}{4} \\ = \frac{(k+1)^2\cdot [k^2 + 4(k+1)]}{4} \\ = \frac{(k+1)^2\cdot [k^2 + 4k+4]}{4} \\ = \frac{(k+1)^2\cdot [(k+2)(k+2)]}{4} \\ = \frac{(k+1)^2\cdot (k+2)^2}{2^2} \\ = \left( \frac{(k+1)\cdot (k+2)}{2} \right)^2 } $$

OpenStudy (anonymous):

holy moly that looks like magic! hahaha I didn't think of doing it like that. For b, could you check to where I got to? It would be really helpful :/

OpenStudy (anonymous):

first wouldnt the next term be (k+1) * (k+1)! ?

OpenStudy (perl):

yes sorry

OpenStudy (perl):

$$ \Large \rm { 1 · 1! + 2 · 2! + . . . + n · n! = (n + 1)! − 1 \\ Basis ~ Case : \\ 1\cdot 1! = (1+1)! -1 ~~\checkmark \\ Inductive ~ Step:\\~\\ \\Assume ~that\\~\\ ~ ~ 1 · 1! + 2 · 2! + . . . + k· k! = (k + 1)! − 1\\\\~\\ \\ Then ~ it ~ follows~\\~\\ \\ 1 · 1! + 2 · 2! + . . . + k· k! + (k+1) \cdot (k+1)! \\= \color{blue}{(k + 1)! − 1} + (k+1) \cdot (k+1)! } $$

OpenStudy (anonymous):

so then how would you reduce that? if your ( k+1) * (k+1) doesn't allow you to do anything else because of the multiplication, no?

OpenStudy (anonymous):

I see.... so that would be the same as the initial problem. Thank you !

OpenStudy (perl):

$$ \Large \rm { 1 · 1! + 2 · 2! + . . . + n · n! = (n + 1)! − 1 \\ Basis ~ Case : \\ 1\cdot 1! = (1+1)! -1 ~~\checkmark \\ Inductive ~ Step:\\~\\ \\Assume ~that\\~\\ ~ ~ 1 · 1! + 2 · 2! + . . . + k· k! = (k + 1)! − 1\\\\~\\ \\ Then ~ it ~ follows~\\~\\ \\ 1 · 1! + 2 · 2! + . . . + k· k! + (k+1) \cdot (k+1)! \\= \color{blue}{(k + 1)! − 1} + (k+1) \cdot (k+1)! \\= (k + 1)![ 1 +(k+1)] − 1 \\= (k + 1)!(k+2) − 1 \\ = (k+2)! - 1 } $$

OpenStudy (perl):

c) is not as easy

OpenStudy (anonymous):

I thought c would be false, as the basic case is false, no?

OpenStudy (perl):

it is true for n> 6

OpenStudy (perl):

thats what you want to prove

OpenStudy (perl):

the basis case would start at n=7

OpenStudy (anonymous):

oh got caught in always inputting 1... you're right. but then you would have something like this: k+1> 6 then 3^(k+1) > (k+1)! right?. The only thing i can think of is using a log function

OpenStudy (anonymous):

or is this super induction?

OpenStudy (perl):

we have to be more clever here

OpenStudy (perl):

$$ \large \rm { Inductive ~step,~ suppose ~that \\ 3k~ < ~k!\\ Then \\ \\ 3(k+1) = 3k + 3 < \color{blue}{k!} + 3 } $$

OpenStudy (perl):

copied it wrong

OpenStudy (anonymous):

Yeah I copied it wrong its 3^n my apologies :/ thats why I didnt know what to say haha

OpenStudy (perl):

$$ \large \rm { Inductive ~step,~ suppose ~that \\ 3^k~ < ~k!\\ we~ want ~ to ~ prove \\ 3^{k+1}~ < ~(k+1)! \\~we ~ need ~ intermediate~ inequalities \\ 3^{k+1}= 3^k \cdot 3 <~~~~~~~~~~~~<~ k!~(k+1)= ~(k+1)! } $$

OpenStudy (perl):

$$ \large \rm { Inductive ~step,~ suppose ~that \\ 3^k~ < ~k!\\ we~ want ~ to ~ prove \\ 3^{k+1}~ < ~(k+1)! \\~we ~ need ~ intermediate~ inequalities \\ 3^{k+1}= 3^k \cdot 3 <~~~\color{blue}{3^k} \color{red}{(k+1)}<~\color{blue}{ k!}~\color{red}{(k+1)}= (k+1)! } $$

OpenStudy (perl):

so this is true if (k+1) > 3 , and it is since k >= 7

OpenStudy (perl):

k + 1 > 3 is equivalent to k > 2

OpenStudy (anonymous):

I don't really see how you got that intermediate inequality, it seems like you pulled it out of nowhere

OpenStudy (perl):

start with the right side (k+1)!

OpenStudy (anonymous):

OH i see, take part of each expression

OpenStudy (anonymous):

then why not take 3*k! as the intermediate

OpenStudy (perl):

the inductive step is (lets read this from right to left) k ! > 3^k Therefore (k+1)! = k! ( k+1) > 3^k * (k+1) , but k+1 > 3 since k >= 7

OpenStudy (perl):

if you can find an inequality that works, you can use it

OpenStudy (perl):

the inductive step is: reading from right to left Assume k ! > 3^k Then (k+1)! = k! ( k+1) > 3^k * (k+1) > 3^k * 3 (since k >= 7 , k+1 > 3 )

OpenStudy (anonymous):

two last things if you don't mind ( im really sorry). How does finding an intermediate inequality proves the initial inequality? secondly, why do you need to highlight k is greater or equal to 7 THEN k+1 > 3. Isn't the second implied in the first part of the statement?

OpenStudy (perl):

because then we proved that (k+1)! > 3^k+1

OpenStudy (perl):

if 10 > 3 > 1 , then 10 > 1 , we can get rid of the intermediate inequality since we are not interested in it in the final result

OpenStudy (perl):

but we need the intermediate inequality to justify it

OpenStudy (anonymous):

Oh I see if you can prove both side of the intermediate inequality, then you can justify the initial statement. been years since ive done that. Thanks a lot for your patience. youre amazing :)

OpenStudy (perl):

right, its like saying if 1 < 3 < 5 < 10 , then 1 < 10

OpenStudy (perl):

so you just have to start at one side and go forward or backwards,

OpenStudy (perl):

one side of 3^(k+1) < (k+1)! you can start at the left hand side. or you can start at the right hand side

OpenStudy (anonymous):

I see, it makes loads of sense. again thank you so so much

OpenStudy (perl):

the best way to do it is to have both sides of the inequality to look at . so start with 3^(k+1) < < (k+1)! Then begin to fill in the space, you want to close this inequality gap with logically true inequalities.

OpenStudy (perl):

3^(k+1) < < (k+1)! ok might as well expand the left and right hand sides 3^(k+1)=3^k * 3 < < k! (k+1) = (k+1)!